a particle moves along a straight line such that its displacement at any time t is given by s=t^3-3t^2+2 m .the displacement when the acceleration becomes zero

Grade:11

1 Answers

Himanshu Rathour
5 years ago
Given that displacement s=t3-3t²+2...... (i) 
We know that v=ds/dt
Differentiating i wrt t
v=3t²-6t
Now a=dv/dt
a=6t-6
putting a=0
t=1
Now, displacement when a=0 i.e. t=1
S=0m